1. Student obtained the value 26.5 for the slope of the fitting straight line on the graph P=const*(1/V) representing the experimental data from the water-based 2-tube apparatus. The experiment was conducted at temperature 22 degree Celsius. What value for the ideal gas constant follows from the obtained slope? The additional necessary data are provided in the text.

2. In the pressure versus temperature experiment, the graph P=const*(T+273.15) of the experimental data (temperature in degree Celsius) was fitted with a straight line, which appeared to have a slope of 345 and an intercept of 93770. What temperature (in degree Celsius) will be defined as the absolute zero by the extrapolation of this fitting line towards zero pressure?
Answer #1

V(1 mol) = (22.414x10^-3)(T)/ 273.15 K = (22.414x10^-3)(299.65 K)/ 273.15 K = 0.0245

n = ? r^2 Lo / V (1 mol) = (? 0.0095^2)(0.917m)/ 0.0245 = 0.0106

R = "graph slope"/nT = 26.5/ (0.0106)(299.65 K) = 8.343

This is the first part. Still working on the second question.